If sin(x) = 3/5 and cos(a) = 7/25 and (x) and (a) are both acute angles find cos((x) + (a)) ALGEBRAICALLY!
Answer by lwsshak3(11628)

If sin(x) = 3/5 and cos(a) = 7/25 and (x) and (a) are both acute angles find cos((x) + (a)) ALGEBRAICALLY!
..
sinx=3/5=.6
cos^2x=1-sin^2x=1-(3/5)^2=1-9/25=16/25
cosx=sqrt(16/25)=4/5=.8
..
cosa=7/25=.28
sin^2a=1-cos^2a=1-(7/25)^2=1-49/625=.92160
sina=sqrt(.92160)=.96
Use addition formula for cos to solve:
cos(s+t)=cos s cos t - sin s sin t
cos(x+a)=cos x cos a-sin x sin a
=.8*.28-.6*.96
=.224-.576
cos(x+a) =-.352